OpenStudy (aroub):

In a ship race Hana completed 20 km downstream in 3 hr. If the return trip took 4 hr 20 min, how fast can she paddle in still water?

hero (hero):

d = rt 20 = 3(r+w) 20 = (13(r+w))/3

OpenStudy (aroub):

Yes, this is what I did and then got a negative answer!

hero (hero):

well, the second equation is supposed to be r - w for upstream

hero (hero):

Let me finish solving

OpenStudy (aroub):

Okay

OpenStudy (aroub):

nvm, I think I did something wrong in the calculations -.-

OpenStudy (anonymous):

Calculated in jiffy, by any chance is it \( \frac{220}{39} km/hr\)?

hero (hero):

FFM, I would swear that your comment is a violation of the code of conduct.

hero (hero):

Posting answers, with no steps and butting in while someone else is trying to help.

OpenStudy (aroub):

I got 95/3

OpenStudy (anonymous):

No need to get all bummed up. I was just testing my mind.

OpenStudy (anonymous):

Here is how I got that : \[ x+y= 20/3 \] \[ x-y= 20\left/\frac{13}{3}\right. \] No solving for \( x\)

OpenStudy (anonymous):

now*

hero (hero):

Yeah, I meant to do it that way

OpenStudy (aroub):

But you'll get x AND y, what do you do?

OpenStudy (anonymous):

x = speed in still water y= speed of the steam/river

hero (hero):

That's why I used r and w

hero (hero):

Did anybody get r = 5.64103 and w = 1.02563

OpenStudy (aroub):

FFM did!

OpenStudy (aroub):

Okay.. Thanks a lot!!